Operation Summary for August 15, 2011 - August 17, 2011 |
|
Monday August 15 The day shift began with the Tevatron (TEV) off for a Kautzky Valve leak repair, with Operations monitoring the Recycler stash (301.7E10) and the Antiproton Source (Pbar) stack (127.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBFT-1010experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, FESS personnel reported at 8:22 AM that they had valved off the Linac ICW to repair a leak, which shut off Linac beam. Operators prepared the Pelletron for access at 8:56 AM. FESS personnel began repairs on a TEV sector E1 LCW leak. Preaccelerator personnel began work on the H- Source at 9:37 AM. Operations resumed sending beam to MiniBooNE at 9:49 AM. MiniBooNE turned on after vacuum and LCW work at 9:50 AM, ran for an hour and then shut off for 4.5 hours to repair their Resistive Wall Monitor. Pelletron personnel reported that they had completed their work at MI-31 at 10:54 AM; however, a breaker trip required another access. Operations put the TEV through a wet squeeze at 1:24 PM. Cryo system personnel reported at 2:37 PM that the TEV F3 wet engine needed at flywheel change; they began repairs. MiniBooNE turned back on about 3:30 PM. The Pelletron suffered from a series of full discharges throughout the day. No beam was transferred to the Recycler during day shift. During the evening shift, Operations put the TEV into shot setup for a 1x1 (1 bunch of protons and 1 bunch of antiprotons) study that ran until 7:10 PM. Operators conducted a series of orbit checks that lasted until 9:20 PM. Operations put the TEV into shot setup, but this was interrupted by BPM problems. Operations established store 8983 at 11:07 PM with an average initial luminosity (AIL) of 187.8E30. Operators monitored the Pelletron, which was in conditioning mode all shift. Tuesday August 16 The midnight shift began with the Pelletron conditioning, with Operations monitoring store 8983 (136.54E30) and the stash (65.1E10), with Pbar stacking (261.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operators attempted to transfer antiprotons to the Recycler, but the transfer was delayed for almost 5 hours by an update to an ACNET code library that broke several ACNET scripts used for transfers. Once an old code library was loaded, operators filled the Recycler to 130E10, which was near to the maximum amount of beam that can be cooled with stochastic cooling only. Operations established beam to T-1010 at 4 AM. The Pelletron returned to normal service at 6:30 AM. During the day shift, operators reported a small problem with the Pelletron at 11:34 AM; it recovered. Operations terminated store 8983 at 12:30 PM, put the TEV into shot setup, and established store 8984 (a 3x3 study store) at 1:38 PM with an AIL of 22.22E30. At 3:38 PM, operators reported that a Booster RF station (BRF1) tripped off and would not reset; it will need a tunnel access to fix. At 3:52 PM, the Booster East Anode Power Supply tripped off with no indication of a cause; Booster RF personnel investigated BRF7. During the evening shift, operators loaded an Energy Conservation Timeline into the Time Line Generator (TLG) at 4:03 PM. Booster RF technicians bypassed BRF7 and Operations resumed sending beam. Booster RF techs brought BRF7 back online at 5:32 PM. Operations terminated study store 8984 at 6:58 PM. Operations disabled beam to T-1010 at 7:12 PM; this ended the experiment’s run. Operations put the TEV into shot setup at 7:12 PM. established store 8985 at 8:17 PM with and AIL of 357.34E30. At 9:21 PM, BRF7 tripped off for the first of three times in two hours; it reset. Wednesday August 17 The midnight shift began with Operations monitoring store 8985 (63.59E30) and the stash (275.4E10), with Pbar stacking (43.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, BRF7 tripped off at 12:43 AM; it reset. Operators tuned the 750 KeV line at 4:21 AM. Operators tuned Booster injection at 6:28 AM. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow TEV personnel to conduct a 3x3 study store, for Booster personnel to conduct studies, and for Booster RF techs to bring BRF1 back online. |
|
More Information Comments and Suggestions |
| last modified 08/17/2011 email Fermilab |
| Security, Privacy, Legal | |